### v2

@@ 5.5 Effect of BrowserGym Features
More is not always better:  As we gradually add more features, the prompt becomes longer and seems to overwhelm the LLM. This is mostly observed on WorkArena where additional features lead to a performance drop. We hypothesise that since the AXTree is already large, adding more features is likely to distract the agent instead of helping it. We also note that in WorkArena, more features imply more prompt truncation with GPT-3.5 and CodeLlama, due to limited context length (Fig. 4).
@@ 6 Conclusion
In this work, we introduced WorkArena, a new benchmark for the evaluation of web agents on tasks inspired by the day-to-day workflow of knowledge workers in the ServiceNow platform. We also introduced BrowserGym, a robust, general-purpose environment for automated web agents, which encompasses an extensive list of features previously proposed in the literature (DOM, AXTree, screenshot, code and high-level action space), as well as novel capabilities such as an interactive chat.
We presented an empirical evaluation of GPT-3.5, GPT-4 and CodeLlama – among the most advanced general-purpose Large Language Models (LLMs) currently available for intruction-following and coding. Specifically, we investigated their generalization performance as web agents in both WorkArena and MiniWoB. Our results validate WorkArena as an unsolved, challenging benchmark that requires advanced reasoning capabilities over long contexts (DOM or AXTree), which seem to emerge only in very large models.
Due to the multi-modal nature of web observations (textual and visual), and the potentially unlimited complexity of web tasks that can be designed (from toy setups like MiniWoB to harder benchmarks like WebArena and WorkArena), we believe that browser-based task automation provides the perfect testbed to evaluate the emergent capabilities of multimodal large language models. In future work, we plan to integrate additional standard benchmarks into BrowserGym, such as WebShop (Yao et al., 2022), WebArena (Zhou et al., 2023), and WebVoyager (He et al., 2024). We also plan to expand WorkArena with tasks covering other, more visual parts of the ServiceNow platform (dashboards, workspaces, application building), and compositional tasks based on common user trajectories inspired by the platform’s certification curricula (e.g., Business Process Analyst).
Both WorkArena and BrowserGym are provided as open-source contributions to the community, and are meant to serve as a catalyst to accelerate both the development of new web agents and their evaluation in terms of capability and potential impact on the real world.
@@ A.1 Tasks
Table 4: List of all tasks available in WorkArena, grouped by category. The number of instances corresponds to the number of instantiations of the parameters of the tasks (e.g., values to input into a specific field). Due to the combinatorial nature of list and form tasks, which resulted in an exceedingly large pool of potential instances, we chose to cap the number of instances at 1,000, selected randomly.
Category Task Name Number of instances
Lists (12 tasks) FilterAssetList 1,000
FilterChangeRequestList 1,000
FilterHardwareList 1,000
FilterIncidentList 1,000
FilterServiceCatalogItemList 1,000
FilterUserList 1,000
SortAssetList 150
SortChangeRequestList 150
SortHardwareList 150
SortIncidentList 150
SortServiceCatalogItemList 150
SortUserList 150
Forms (5 tasks) CreateChangeRequest 1,000
CreateIncident 1,000
CreateHardwareAsset 1,000
CreateProblem 1,000
CreateUser 1,000
Knowledge Bases (1 task) KnowledgeBaseSearch 1,000
Service Catalogs (9 tasks) OrderDeveloperLaptopMac 1,000
OrderIpadMini 80
OrderIpadPro 60
OrderSalesLaptop 1,000
OrderStandardLaptop 1,000
OrderAppleWatch 10
OrderAppleMacBookPro15 10
OrderDevelopmentLaptopPC 40
OrderLoanerLaptop 350
Menus (2 tasks) AllMenu 1,000
Impersonation 600
Total 18,050
@@ A.2 Task User Interface Examples
In this section, we provide an example of the typical user interface encountered for each category of task (Figs. 5, 6, 7 and 8). We omit “form” tasks, as such an example has already been presented in Fig. 3.
@@ A.3 Knowledge Base Tasks – Additional Details
The knowledge base included in WorkArena consists of 100 articles generated using GPT-4 (OpenAI, 2023). To achieve this, we start from a list of 100 facts, which are each composed of an item and a value. Tab. 5 shows a few examples. Then, for each fact, we use GPT-4 to produce an article in HTML format and make sure that the exact string “the {fact} is {item}” is included in the article. An example article is shown in Fig. 9.
@@
For each fact, i.e., (item, value) pair, we produce a list of questions that ask about item and whose answers are exactly value. We achieve this by prompting GPT-4 with the initial question “What is {item}?” and ask it to produce 10 alternative wordings for the question, as well as formatting instructions, to ensure that the answer is exactly value. Then, we prompt GPT-3.5 with the generated article and each question, ensuring that every single one is answered correctly. If the model fails to answer a question, we ask GPT-4 to improve it and we repeat the process. Note that we use GPT-3.5 to answer the questions to avoid the pitfall where GPT-4 would cater to itself, producing ambiguous questions that it somehow succeeds in answering correctly. Example questions and formatting instructions are shown in Tab. 6.
@@
Despite the precise formatting instructions included with each question, we allow for slight variations in formatting and wording by verifying if the answer produced by the agent is within a set of acceptable answers. To produce such alternative answers, we provide the expected value to GPT-4 and ask it to produce 10 alternative formats. We include multiple examples in the prompt and inspect the results to ensure coherence. An example is shown in Tab. 7.
@@ B.2 MiniWoB
Appendix C Analyzing CodeLLAMA’s Shortcomings in WorkArena
This section delves into specific failure modes exhibited by CodeLLAMA within the WorkArena environment, with orange text emphasizing key aspects of these shortcomings.
C.1 Mimicking Actions from the Action Space Description
In the scenario presented below, the agent navigates the BID (high-level) action space. Rather than selecting an actionable step aligned with the objective and accessibility structure, it opts for an action directly from the exemplars provided in the action space description.
Prompt
1# System message
2
3You are an agent trying to solve a web task based on the content of the page and a user instructions. You can interact with the page and explore. Each time you submit an action it will be sent to the browser and you will receive a new page.
4
5# Instructions
6
7Review the current state of the page and all other information to find the best possible next action to accomplish your goal. You are in a benchmark and your answer will be interpreted and executed by a program, make sure to follow the formatting instructions.
8
9## Goal:
10
11Navigate to the "Live Workflows > History" module of the "Workflow" application.
12
13# Observation of current step:
14
15## AXTree:
16...
17
18# History of interaction with the task:
19
20# Action space:
21
2214 different types of actions are available.
23
24select_option(bid: str, options: str | list[str])
25 ...
26
27dblclick(bid: str, button: Literal[’left’, ’middle’, ’right’] = ’left’, modifiers: list[typing.Literal[’Alt’, ’Control’, ’Meta’, ’Shift’]] = [])
28 ...
29
30focus(bid: str)
31 ...
32
33hover(bid: str)
34 ...
35
36scroll(delta_x: float, delta_y: float)
37 ...
38
39send_msg_to_user(text: str)
40 ...
41
42clear(bid: str)
43 ...
44
45noop()
46 ...
47
48press(bid: str, key_comb: str)
49 ...
50
51click(bid: str, button: Literal[’left’, ’middle’, ’right’] = ’left’, modifiers: list[typing.Literal[’Alt’, ’Control’, ’Meta’, ’Shift’]] = [])
52 Description: Click an element.
53 Examples:
54 click(’51’)
55 ...
56
57drag_and_drop(from_bid: str, to_bid: str)
58 ...
59
60fill(bid: str, value: str)
61 ...
62
63Only a single action can be provided at once. Example:
64press(’88’, ’Backspace’)
65Multiple actions are meant to be executed sequentially without any feedback from the page.
66Dont execute multiple actions at once if you need feedback from the page.
67
68# Abstract Example
69
70Here is an abstract version of the answer with description of the content of each tag. Make sure you follow this structure, but replace the content with your answer:
71
72*action*
73One single action to be executed. You can only use one action at a time.
74*action*
75
76# Concrete Example
77
78Here is a concrete example of how to format your answer.
79Make sure to follow the template with proper tags:
80
81*action*
82fill(’237’, ’example value’)
83*action*
Agent’s Response
1*action*
2click(’51’)
3*action*
As illustrated, CodeLLAMA directly replicated the example action listed on line 61.
C.2 Inadequate Utilization of the BID API
This example highlights CodeLLAMA’s errors in action selection and coding syntax. Initially, it opts for an action outside the set of permitted actions. Subsequently, it commits a syntax error in the code.
Prompt
1
2# System Message
3...
4
5# Instructions
6...
7
8## Goal:
9Sort the "assets" list by the following fields:
10 - State (ascending)
11
12# Observation of current step:
13
14## AXTree:
15...
16
17# History of interaction with the task:
18
19# Action space:
20
21
2214 different types of actions are available.
23
24select_option(bid: str, options: str | list[str])
25 ...
26
27dblclick(bid: str, button: Literal[’left’, ’middle’, ’right’] = ’left’, modifiers: list[typing.Literal[’Alt’, ’Control’, ’Meta’, ’Shift’]] = [])
28 ...
29
30focus(bid: str)
31 ...
32
33hover(bid: str)
34 ...
35
36scroll(delta_x: float, delta_y: float)
37 ...
38
39send_msg_to_user(text: str)
40 ...
41
42clear(bid: str)
43 ...
44
45noop()
46 ...
47
48press(bid: str, key_comb: str)
49 ...
50
51click(bid: str, button: Literal[’left’, ’middle’, ’right’] = ’left’, modifiers: list[typing.Literal[’Alt’, ’Control’, ’Meta’, ’Shift’]] = [])
52 ...
53
54drag_and_drop(from_bid: str, to_bid: str)
55 ...
56
57fill(bid: str, value: str)
58 ...
59
60
61Note:
62...
63
64# Abstract Example
65...
66
67# Concrete Example
68...
Agent’s Response
1*action*
2sort(’208-192’, ’State’, ’ascending’)
3*action*
Enviroment Feedback
1Error while parsing action: invalid action type sort.
2Make sure your answer is restricted to the allowed actions.
Agent’s Response
1I apologize for the mistake. You are correct, the action type "sort" is not allowed.
2
3Here is a revised version of the answer that uses a valid action type:
4
5*action*
6fill(’208-192’, ’State’, ’ascending’)
7*action*
8
9I hope this revised answer meets your requirements. Please let me know if you have any further questions or concerns.
Enviroment Feedback
1TypeError: fill() takes 2 positional arguments but 3 were given
As demonstrated, the agent initially selected the sort function, which does not exist. Following this, it encountered a syntax error while attempting to use the fill function.
C.3 Incorrect Application of the Python API
Subsequently, CodeLLAMA misapplies the Python API in an attempt to generate Playwright code.
Prompt
1# System message
2...
3
4# Instructions
5...
6
7## Goal:
8Navigate to the "Knowledge" module of the "Self-Service" application.
9
10# Observation of current step:
11
12## AXTree:
13...
14
15# History of interaction with the task:
16
17# Action space:
18
19...
20The browser context is in page.context, and all open webpages (tabs)
21are in page.context.pages. Here is an example of a valid action:
22
23frame = page.frame_locator(".result-frame")
24button = frame.get_by_text("Submit")
25button.click()
26
27Here is another example:
28
29frame = page.get_by_test_id("35").frame_locator(":scope")
30frame.get_by_test_id("35-776").click()
31
32...
33
34# Abstract Example
35
36Here is an abstract version of the answer with description of the content of
37each tag. Make sure you follow this structure, but replace the content with your
38answer:
39
40*action*
41One single bloc of Python code. Do not include any explanation, only valid Python code.
42*action*
43
44# Concrete Example
45
46Here is a concrete example of how to format your answer.
47Make sure to follow the template with proper tags:
48
49*action*
50frame = page.get_by_test_id("23").frame_locator(":scope")
51frame = frame.get_by_test_id("23-557").frame_locator(":scope")
52frame.get_by_test_id("23-557-2").fill("Hello world!")
53frame.get_by_test_id("23-557-3").click()
54*action*
Agent Response
1*action*
2page.get_by_test_id("Knowledge").click()
3*action*
As one witnesses, CodeLLAMA confuses the get_by_test_id and the get_by_text functions.
C.4 Struggling with iFrame Complexity
In the final example, CodeLLAMA incorrectly navigates the complexities arising from handling iFrames within Playwright.
Prompt
1# System message
2...
3
4# Instructions
5...
6
7## Goal:
8Go to the hardware store and order 4 "Apple Watch"
9
10# Observation of current step:
11
12## AXTree:
13...
14
15# History of interaction with the task:
16
17# Action space:
18...
19
20Note that Playwrights get_by_test_id() method is configured to use the bid attribute to locate HTML elements, instead of the default data-testid. Also, Playwrights locators can not traverse iframes, so you have to locate parent iframes first in order to locate an element in an iframe. The bid attribute contains all the information required to recursively locate an element. For example, an element with bid="23-557-2" can be retrieved as follows:
21
22frame = page.get_by_test_id("23").frame_locator(":scope")
23frame = frame.get_by_test_id("23-557").frame_locator(":scope")
24elem = frame.get_by_test_id("23-557-2")
25...
26
27# Abstract Example
28
29Here is an abstract version of the answer with description of the content of each tag. Make sure you follow this structure, but replace the content with your answer:
30
31*memory*
32Write down anything you need to remember for next steps. You will be presented with the list of previous memories and past actions. Some tasks require to remember hints from previous steps in order to solve it.
33*memory*
34
35*action*
36One single bloc of Python code. Do not include any explanation, only valid Python code.
37*action*
38
39# Concrete Example
40
41Here is a concrete example of how to format your answer.
42Make sure to follow the template with proper tags:
43
44*memory*
45I clicked on bid 32 to activate tab 2. The accessibility tree should mention focusable for elements of the form at next step.
46*memory*
47
48*action*
49frame = page.get_by_test_id("23").frame_locator(":scope")
50frame = frame.get_by_test_id("23-557").frame_locator(":scope")
51frame.get_by_test_id("23-557-2").fill("Hello world!")
52frame.get_by_test_id("23-557-3").click()
53*action*
Agent response
1*memory*
2
3* I am currently on the ServiceNow Catalog page.
4* I have not yet interacted with the page.
5* I am trying to order 4 "Apple Watch" items.
6
7*memory*
8
9*action*
10
11# Find the search bar and enter "Apple Watch"
12search_bar = page.get_by_test_id("212-88")
13search_bar.fill("Apple Watch")
14
15# Wait for the search results to load
16page.wait_for_load_state("networkidle")
17
18# Find the "Apple Watch" result and click on it
19...
20
21*action*
As illustrated, CodeLLAMA overlooks the limitation that Playwright cannot traverse iframes directly. It fails to acknowledge the need to first locate parent iframes before attempting to locate an element within an iframe.

### v5

@@ 5.4 Ablation Study
More is not always better:  As we gradually add more features, the prompt becomes longer and seems to overwhelm the LLM. This is mostly observed with GPT-3.5 and Llama3, the weaker LLMs, where seemingly harmless features such as a long_description or individual_examples in the action space description actually degrade performance. We hypothesise that since the AXTree is already large, adding too much non-crucial features is likely to distract the agent instead of helping it. We also note that in WorkArena, more features imply more prompt truncation for GPT-3.5 and Llama3 due to their limited context length (Figure 4).
Memory can hurt:  The use_think_history flag provides an interesting capacity to our agents, that of remembering their thoughts from previous time steps. We observed in practice that this features works, in the sense that the agents do re-use information about previous pages. However, such a feature is mostly useful when the task requires some form of memory, such as navigating a first page to find some information (e.g., an address), and then going to another page (e.g., Google maps) to type this information. On both WorkArena and MiniWoB the vast majority of tasks does not require navigation, and all the necessary information is always present on the current page. What we observe instead, is that agents with this feature activated tend to stick to decisions decided in early time steps, even erroneous ones, and are somehow less prone to self-correction. This is mostly observed with the GPT agents, Tables 3 and 4.
Deceiving multimodal performance:  We report in Table 2 a complimentary experiment with GPT-4o-V, a vision-augmented GPT-4o model, to assess the capacity of multi-modal models at utilizing the screenshot of the current web page for solving web tasks. We augment the screenshot with Set-of-Mark (Yang et al., 2023) as proposed in WebVoyager (He et al., 2024), and we keep the rest of the flags identical to the GPT-4o best performing agent. The results are somewhat deceiving, with very minor performance improvements on MiniWoB and WebArena. Similarly poor multimodal performances are reported by Xie et al. (2024), which could be explained by the lack of screen-related training data in current vision-language models. This leaves room for future development in this area.
2D features do not always help:  Several MiniWoB tasks require precise 2D understanding of the UI and interaction with $x,y$ coordinates. While we observe a notable performance gain on this benchmark when giving GPT-4o additional coordinate features and actions (Table 3), this does not translate to WorkArena where the tasks seem best solved using bid only. This seems to indicate that, while 2D actions are clearly required in specific cases (i.e., drawing on a web whiteboard), most web interactions actually do not require it, which can greatly simplify the agent design.
Figure 4: Comparative size analysis of different page observation modalities: HTML (left) and accessibility tree (right) across MiniWoB, WorkArena and WebArena.
@@ 6 Conclusion
In this work, we introduced WorkArena, a new benchmark for the evaluation of web agents on tasks inspired by the day-to-day workflow of knowledge workers in the ServiceNow platform. WorkArena underscores the challenge of navigating real-world enterprise websites, notably dealing with large observations and complex, non-standard interfaces.
We also introduced BrowserGym, a robust, general-purpose environment for automated web agents, which encompasses an extensive list of features previously proposed in the literature (HTML, AXTree, screenshot, set-of-mark, code and high-level action space), as well as novel capabilities such as an interactive chat. BrowserGym allows for the flawless evaluation of web agents on multiple benchmarks within the same unified framework, and currently supports MiniWoB, WebArena and WorkArena. Both contributions are open source and meant to serve as catalysts to accelerate the development of new web agents and their evaluation in terms of capability and potential impact on the real world.
We presented an empirical evaluation of GPT-3.5, GPT-4o and Llama3 – among the most advanced general-purpose Large Language Models (LLMs) currently available for intruction-following and coding. Specifically, we investigated their generalization performance as web agents in MiniWoB, WebArena and WorkArena. Our results validate WorkArena as an unsolved, challenging benchmark that requires advanced reasoning capabilities over long contexts (HTML or AXTree), which seem to emerge only in very large models.
In future work we plan to integrate additional standard benchmarks into BrowserGym, such as WebShop (Yao et al., 2022) and WebVoyager (He et al., 2024). We also plan to expand WorkArena significantly by creating compositional tasks built upon the present benchmark, to obtain realistic complex workflows that require additional skills like retrieval, memorization, visual perception, and advanced reasoning.
Due to the multi-modal nature of web observations (textual and visual), and the potentially unlimited complexity of web tasks that can be designed (from toy setups like MiniWoB to harder benchmarks like WebArena and WorkArena), we believe that browser-based task automation provides the perfect testbed to evaluate the emergent capabilities of multimodal large language models. We hope that the present work will stimulate further progress in the community.
@@ A.1 Tasks
Table 6: List of all tasks available in WorkArena, grouped by category. Instances: The number of instances corresponds to the number of instantiations of the parameters of the tasks (e.g., values to input into a specific field). Due to the combinatorial nature of list and form tasks, which resulted in an exceedingly large pool of potential instances, we chose to cap the number of instances at 1,000, selected randomly. Oracle Actions: The number of Playwright actions required by the oracle functions (see § 4) to solve the tasks. This is indicative of task complexity, but we emphasize that human-coded oracle functions, while correct, might not correspond to optimal solutions. Results are averaged over 10 randomly sampled instances and shown along with the standard deviation.
Category Task Name Instances Oracle Actions
Lists (12 tasks) FilterAssetList 1,000 17.3 $\pm$ 6.5
FilterChangeRequestList 1,000 18.7 $\pm$ 4.7
FilterHardwareList 1,000 18.4 $\pm$ 5.6
FilterIncidentList 1,000 16.2 $\pm$ 4.2
FilterServiceCatalogItemList 1,000 19.9 $\pm$ 5.9
FilterUserList 1,000 12.7 $\pm$ 3.2
SortAssetList 150 7.4 $\pm$ 2.3
SortChangeRequestList 150 7.7 $\pm$ 1.6
SortHardwareList 150 8 $\pm$ 2.3
SortIncidentList 150 8 $\pm$ 2.7
SortServiceCatalogItemList 150 8.3 $\pm$ 2.5
SortUserList 150 7.7 $\pm$ 2.1
Forms (5 tasks) CreateChangeRequest 1,000 21.5 $\pm$ 6.2
CreateIncident 1,000 23 $\pm$ 7.9
CreateHardwareAsset 1,000 47.1 $\pm$ 10.9
CreateProblem 1,000 10 $\pm$ 3.4
CreateUser 1,000 17.9 $\pm$ 5.2
Knowledge Bases (1 task) KnowledgeBaseSearch 1,000 4.0 $\pm$ 0.0
Service Catalogs (9 tasks) OrderDeveloperLaptopMac 1,000 8.7 $\pm$ 0.9
OrderIpadMini 80 6.0 $\pm$ 0.0
OrderIpadPro 60 6.0 $\pm$ 0.0
OrderSalesLaptop 1,000 9.0 $\pm$ 0.8
OrderStandardLaptop 1,000 8.0 $\pm$ 0.6
OrderAppleWatch 10 4.0 $\pm$ 0.0
OrderAppleMacBookPro15 10 4.0 $\pm$ 0.0
OrderDevelopmentLaptopPC 40 6.0 $\pm$ 0.0
OrderLoanerLaptop 350 8.0 $\pm$ 0.0
Menus (2 tasks) AllMenu 1,000 3.0 $\pm$ 0.0
Impersonation 600 7.0 $\pm$ 0.0
Dashboards (4 tasks) SingleChartValueRetrieval 1000 1.0 $\pm$ 0.0
SingleChartMinMaxRetrieval 346 1.0 $\pm$ 0.0
MultiChartValueRetrieval 444 2.0 $\pm$ 0.0
MultiChartMinMaxRetrieval 72 2.0 $\pm$ 0.0
Total (33 tasks) 19,912
@@ A.2 Task User Interface Examples
In this section, we provide an example of the typical user interface encountered for each category of task (Figures 5, 6, 7, 9 and 8). We omit “form” tasks, as such an example has already been presented in Figure 3.
@@
Figure 8: Example “MultiChartValueRetrieval” Task – The goal is given to the agent in natural language. As can be seen, the goal is designed to be very explicit, clearly stating which question must be answered and the expected format. Here, the agent must scan the entire dashboard to locate the relevant plot  1. It must then search for the requested label and retrieve its value in the desired format (count or percentage)  2. Finally, it must return this information to the user via the chat box for validation  3.
@@ A.3 Knowledge Base Tasks – Additional Details
The knowledge base included in WorkArena consists of 100 articles generated using GPT-4 (OpenAI, 2023). To achieve this, we start from a list of 100 facts, which are each composed of an item and a value. Table 7 shows a few examples. Then, for each fact, we use GPT-4 to produce an article in HTML format and make sure that the exact string “the {fact} is {item}” is included in the article. An example article is shown in Figure 10.
@@
For each fact, i.e., (item, value) pair, we produce a list of questions that ask about item and whose answers are exactly value. We achieve this by prompting GPT-4 with the initial question “What is {item}?” and ask it to produce 10 alternative wordings for the question, as well as formatting instructions, to ensure that the answer is exactly value. Then, we prompt GPT-3.5 with the generated article and each question, ensuring that every single one is answered correctly. If the model fails to answer a question, we ask GPT-4 to improve it and we repeat the process. Note that we use GPT-3.5 to answer the questions to avoid the pitfall where GPT-4 would cater to itself, producing ambiguous questions that it somehow succeeds in answering correctly. Example questions and formatting instructions are shown in Table 8.
@@
Despite the precise formatting instructions included with each question, we allow for slight variations in formatting and wording by verifying if the answer produced by the agent is within a set of acceptable answers. To produce such alternative answers, we provide the expected value to GPT-4 and ask it to produce 10 alternative formats. We include multiple examples in the prompt and inspect the results to ensure coherence. An example is shown in Table 9.